QUESTION 9

A three- stage amplifier circuit has first stage gain of 45dB, second stage gain of 50dB and 

third stage gain of -5dB. What is the overall gain? If input to the first stage is 0.1mV, what is 

the output of the final stage?



1
Expert's answer
2021-10-18T03:44:56-0400

Overall gain is

"=45+50-5=90dB\\
10\log{\frac{V_o}{V_i}}=90\\
V_i=0.1mV\\
V_o=10^9\times0.1mV=10^8mV"